3. A pillar in the shape of a cone trunk is lying on the ground.

It is 5m long and weighs 600kg. Its center of gravity is located 2m away from the base. What forces must we exert to raise the pillar once from one end and once from the other end?


1
Expert's answer
2021-03-10T17:22:07-0500

1)


"(600)(9.8)(2)=(5-2)F\\F=3920\ N"

2)


"(600)(9.8)(5-2)=(2)F'\\F'=8820\ N"